Assume that your parents wanted to have $35,000 saved for college by your 17th birthday and they started saving on your first birthday. They saved the same amount each year on your birthday and earned 8% per year on their investments. a. How much would they have to save each year to reach their goal? b. If they think you will take five years instead of four to graduate and decide to have $10,000 saved just in case, how much more would they have to save each year to reach their new goal?
Answer #1
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
35000= Cash Flow*(((1+ 8/100)^17-1)/(8/100))
Cash Flow = 1037.03

Part 2

FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
45000= Cash Flow*(((1+ 8/100)^17-1)/(8/100))
Cash Flow = 1333.32

additional amount = 1333.32-1037.03

=

296.29